OpenStudy (anonymous):

Help Please: If A = x^2 and dx/dt =3 when x=10, find dA/dt.

jimthompson5910 (jim_thompson5910):

A = x^2 dA/dt = d/dt[ x^2 ] dA/dt = 2x * d/dt [ x ] dA/dt = 2x * dx/dt I'll let you finish

OpenStudy (anonymous):

can you please elaborate it little bit more this is a new material for you just little bit more if its possible for u?

jimthompson5910 (jim_thompson5910):

well x = 10 (given) and dx/dt = 3

jimthompson5910 (jim_thompson5910):

also given

jimthompson5910 (jim_thompson5910):

plug these in and evaluate

sam (.sam.):

u can use chain rule as well \[\frac{dA}{dt}=\frac{dA}{dx} \frac{dx}{dt}\]

jimthompson5910 (jim_thompson5910):

yes you could do that too and Sam's way is useful for doing partial derivatives later down the road

OpenStudy (anonymous):

ok thanks